### v1

@@ 3.2.2 On solving LPECs
The practical success of the proposed method depends on the efficient solution of LPECs. There exists several tailored LPEC methods Fang2010 ; Hu2012 ; JaraMoroni2020 ; JaraMoroni2018 , but their implementations are not readily available. Since LPECs are a special case of MPECs, they can also be solved with any of the regularization and penalty-based methods mentioned in Section 2.2. However, it may be non-trivial to deal with cases where these methods converge to spurious solutions and fail to verify B-stationarity.
A practical and straightforward approach to solving LPECs (13) is to reformulate them as an equivalent mixed-integer linear program (MILP), which can be expressed as follows Hu2012 :
@@
$0\leq x_{1,i}\!+\!d_{1,i}\!\leq\!yM,$ $\forall i\in\{1,\ldots,m\},$ (21c)
$0\leq x_{2,i}\!+\!d_{2,i}\!\leq\!(e-y)M,$ $\forall i\in\{1,\ldots,m\},$ (21d)
$\|d\|_{\infty}\leq\rho,$ (21e)
@@
where $e$ is a vector of ones with appropriate dimensions, and $M$ is a sufficiently large positive constant. The constraints (21c) and (21d) correspond to the big-M reformulation of the complementarity constraints in (13c). The binary variables $y\in\{0,1\}^{m}$ effectively select one of the branches of the L-shaped set introduced by the complementarity constraints.
The MILP (21) can be solved with state-of-the-art commercial and open-source solvers such as Gurobi Gurobi or HiGHS Huangfu2018 . In general, due to the exponential complexity of MILPs, this can be computationally expensive. Moreover, big-M formulations can be inefficient as they can lead to rather loose relaxations Hu2008 . In the special case of (21), we argue that the MILP can usually be solved very efficiently in practice for several reasons.
@@
Observe that the constraint (21e) makes the feasible set of the MILP compact. Moreover, for feasible points $x^{k}$, the trust region radius $\rho$ can be arbitrarily small, leading to tight relaxations of the MILP. For small $\rho$, the trust region constraint (21e) always dominates the bounds on $d$ implied by the big-M constraints (21c) and (21d). Therefore, a large $M$ does not lead to loose relaxations of the MILP. Thus, $M$ only needs to be large enough to ensure the feasibility of  (21c)-(21d) and that is not tighter than (21e). One choice is $M=\max((x_{1}^{k},x_{2}^{k}))+\rho$. Furthermore, given a feasible complementarity pair, e.g. from the previous BNLP or LPEC solution, we can construct a feasible binary vector $y$, leading to a good upper bound.
All this together typically leads to faster solution times in practice. As it will be shown in Section 5, the cumulative time spent solving MILPs (21) is only a small fraction of the overall solution time. Moreover, it will also be seen that the MILP formulations for LPECs lead to much smaller computation times than the corresponding regularization-based methods.
We mention also the special but not so uncommon case when a point $x^{*}\in\Omega$ is an S-stationary point. In this case, the MILP is especially easy to solve, because the solution of the relaxed MILP – which is an LP obtained by setting $y\in[0,1]^{m}$ instead of $y\in\{0,1\}^{m}$ – already solves the MILP. This means that the verification of a B-stationary point in this case reduces to a single LP solve. The next lemma formalizes this assertion.
@@
Let $x^{*}\in\Omega$ be a S-stationary point of the MPEC (1). For a sufficiently small $\rho>0$, a global minimizer of the relaxed MILP (21) is $d=0$, and $y\in[0,1]^{m}$ such that (21c) and (21d) hold. Moreover, any $y\in\{0,1\}^{m}$ that satisfies (21c) and (21d) and $d=0$ are a global minimizer of the MILP (21).
Proof. It follows from Theorem 2.1 that if $x^{*}\in\Omega$ is S-stationary, it is also B-stationary, hence $d=0$ is a local minimizer of the ${\mathrm{LPEC}}(x^{*},\rho)$ in (13). For a sufficiently small $\rho>0$, $d=0$ is also a global minimizer of ${\mathrm{LPEC}}(x^{*},\rho)$. Now, since $x^{*}$ is S-stationary, it is also a stationary point of the so-called relaxed NLP, which is obtained by replacing (9e) in the TNLP by $x_{1,i},x_{2,i}\geq 0,i\in{\mathcal{I}}_{00}(x^{*})$ Scheel2000 . This also means that the LP obtained by linearizing all functions of the relaxed NLP has $d=0$ as its global minimizer. Adding a trust region constraint with $\rho>0$ on the step $d$ does not change the solution of this LP, which reads as:
s.t. $c(x^{k})+\nabla c(x^{k})^{\top}d\geq 0,$ (22b)
$x_{1,i}+d_{1,i}=0,\ x_{2,i}+d_{2,i}\geq 0,\;$ $\forall i\in\mathcal{I}_{0+}(x^{*}),$ (22c)
$x_{1,i}+d_{1,i}\geq 0,\ x_{2,i}+d_{2,i}=0,\;$ $\forall i\in\mathcal{I}_{+0}(x^{*}),$ (22d)
$x_{1,i}+d_{1,i}\geq 0,\ x_{2,i}+d_{2,i}\geq 0,\;$ $\forall i\in\mathcal{I}_{00}(x^{*}),$ (22e)
$\|d\|_{\infty}\leq\rho$ (22f)
Now by applying similar reasoning as in the proof of Lemma 3.1, it can be seen that by relaxing $y\in\{0,1\}^{m}$ to $y\in[0,1]^{m}$ in  (21), and for a sufficiently small $\rho>0$, the feasible sets of relaxed MILP  (21) and the LP (22) coincide for all $d\in{\mathbb{R}}^{n}$ components. Since $y$ does not enter the objective, $d=0$ is also a minimizer of the relaxed MILP. Finally, since the ${\mathrm{LPEC}}(x^{*},\rho)$ in (13) has only one global minimum, we conclude that $d=0$ and any $y$ that satisfies (21c) and (21d) are a global minimizer of the equivalent MILP (21). A vector $y\in\{0,1\}^{m}$ is trivially obtained from the evaluation of (21c) and (21d) with $d=0$ and $x^{*}$. ∎
Once the MILP (21) is solved, if $d\neq 0$ we need to identify the active sets ${\mathcal{I}}_{1}(x^{k}+d)$ and ${\mathcal{I}}_{2}(x^{k}+d)$, which requires some partition of ${\mathcal{I}}_{00}(x^{k}+d)$. There is some ambiguity here, and in our implementation, we use the MILP solution and define
${\mathcal{I}}_{1}(x^{k}+d)=\{i\mid y_{i}=0\},\ {\mathcal{I}}_{2}(x^{k}+d)=\{i\mid y_{i}=1\}.$
We highlight that MILP solvers usually solve (21) to global optimality. For the presented method, however, it is sufficient that the LPEC solver either verifies $d=0$ as a local optimizer or finds a descent direction $d$ that predicts an active set and a corresponding BNLP with a lower objective value.
In practice, the global solution may be on an edge of the feasible set, where it is possible to have $\nabla f(x^{k})^{\top}d=0$ with $d\neq 0$, cf. Fig 3a for an example. In such cases, we consistently select $d=0$.
Remark 3.1 (On Algorithm 1 and mixed-integer algorithms)
Many mixed-integer nonlinear programming (MINLP) methods Belotti2013 , solve a sequence of NLPs and MILPs, just like Algorithm 1. Therefore, we highlight their major conceptual differences. Using the big-M reformulation, any MPEC (1) can be reformulated into a MINLP. Most MINLP methods converge to the global optimum when the relaxed integer problem is a convex problem, Algorithm 1 does not. This can be seen in the MPEC example in (20). It can be verified that the equivalent MINLP, obtained via the big-M reformulation, after relaxing the binary variable from $y\in\{0,1\}^{m}$ to $y\in[0,1]^{m}$ results in a convex quadratic program. We have seen in Example 3.3, that the Algorithm 1 does not always converge to the global optimum of the MPEC, but only to a B-stationary point. In summary, Algorithm 1 is not a MINLP algorithm, and the MILP solvers are just a tool to solve LPECs.
3.3 Phase I: Finding a feasible branch NLP
In this section, we introduce two algorithms to find an initial consistent BNLP and its corresponding solution $x^{0}$, or to certify (local) infeasibility of (1) if this is impossible. The first algorithm solves a relaxed NLP, and the second explicitly minimizes the constraint violation subject to the complementarity constraint (1c).
3.3.1 Regularization-based approach
Even though regularization-based methods may not always find B-stationary points, they can be efficient in finding (almost) feasible points. In this section, we present a Phase I method that combines solving LPECs with regularization-based methods described in Section 2.2. In our exposition, we focus on Scholtes’ global relaxation method, cf. Section 2.2. We compute $x^{*}(\tau)$, a stationarity point of Reg($\tau$), which is not necessarily feasible for the MPEC (1). Next, an LPEC is constructed at $x^{*}(\tau)$ (for some $\tau>0$), and the solution of the LPEC is then used to find an active set corresponding to a feasible BNLP. Such an approach is motivated by Lemma 3.2, and the following straightforward fact.
Lemma 3.4
Consider the MPEC (1). Let $c$ be an affine function, $x\in{\mathbb{R}}^{n}$, and $a\in{\mathbb{R}}^{n}$ any nonzero vector. Then, for a sufficiently large $\rho$ and for any local minimizer $d$ of the LPEC:
$\underset{d\in{\mathbb{R}}^{n}}{\mathrm{min}}\;\quad$ $a^{\top}d$ (23a)
s.t. $c(x)+\nabla c(x)^{\top}d\geq 0,$ (23b)
$0\leq x_{1,i}+d_{1,i}\perp x_{2,i}+d_{2,i}\geq 0,$ $\forall i\in\{1,\ldots,m\},$ (23c)
$\|d\|_{\infty}\leq\rho,$ (23d)
it follows, that $x+d\in\Omega$. Additionally, if this LPEC is infeasible for all sufficiently large $\rho>0$, then the MPEC (1) is also infeasible.
This lemma trivially follows from the fact that for $x\in\Omega$ the LPEC and the MPEC have the same feasible set, hence an LPEC with any objective can identify a feasible BNLP. The trust-region radius $\rho>0$ has just to be large enough such that the complementarity constraints can be satisfied. In our implementation, for the objective, we take $a=\nabla f(x)$ (i.e., (23) reduces to (13)), but other choices such as minimizing the distance to the feasible set are also possible.
Now, if the function $c$ is nonlinear, and we have an almost feasible point $x$, we expect that the LPEC (23) can identify a feasible BNLP. We prove below in Theorem 4.1 that this is indeed the case.
On the other hand, certifying that a general nonconvex optimization problem is infeasible is difficult. Therefore, most NLP methods detect so-called local infeasibility, which is defined as follows. If the globalization strategy of an NLP method cannot make sufficient progress toward finding a feasible point, then an auxiliary NLP – which usually minimizes the constraint violation – is solved Fletcher2002 ; Waechter2006 . If a local minimizer (or stationary point) of this auxiliary NLP is not a feasible point of the original NLP, then the original problem is declared to be locally infeasible.
We highlight a fact about infeasibility detection with the regularization-based approach.
Lemma 3.5
@@ 3.3.1 Regularization-based approach
Regard the MPEC (1) and the corresponding $\mathrm{Reg}(\tau)$ in  Reg($\tau$) for some $\tau>0$. If $\mathrm{Reg}(\tau)$ is locally infeasible then the same holds for the MPEC (1).
@@
On the other hand, if Reg($\tau$) is feasible for all $\tau$, then also the MPEC (1) is feasible as well. In a practical implementation, we compute a finite sequence of $x^{*}(\tau^{k})$ for $k=0,\ldots,N^{\tau}$, and solve the $\mathrm{LPEC}(x^{*}(\tau^{k}),\rho)$. We expect that if $\tau^{k}$ is sufficiently small, then the feasible set of $\mathrm{LPEC}(x^{*}(\tau^{k}),\rho)$ is locally a good approximation of the original feasible set. It is not known a priori how small $\tau^{k}$ and how large $\rho$ must be for a feasible BNLP to be identified. The following example illustrates how an LPEC can fail to identify a feasible BNLP if $\tau$ is not sufficiently small.
@@
$\underset{x\in{\mathbb{R}}^{2}}{\mathrm{min}}\;\quad$ $-2x_{1}+x_{2}$
s.t. $-x_{1}-(x_{2}-a)^{2}+1\geq 0$
$0\leq x_{1}\perp x_{2}\geq 0$
with $a>1$. In the example, we set $a=1.1$. If $\tau$ is not small enough, then the LPEC (13) has feasible branches, which are infeasible for the original MPEC. Fig. 4 illustrates two sample solutions for $\mathrm{Reg}(\tau)$ and the corresponding $\mathrm{LPEC}(x^{*}(\tau),\rho)$. In this example, the smaller the parameter $a$ is the smaller $\tau$ must become.
@@
Figure 4: Illustrating of Example 3.4, where the LPEC (13) can correctly and wrongly identify a feasible BNLP depending on the linearization point $x^{*}(\tau)$.

### v2

@@ 3.2.2 On solving LPECs
LPECs are nonconvex and combinatorial problems, and the effectiveness of our proposed method relies on solving them efficiently. In practice, our computational results indicate that although the worst-case complexity of LPECs may be exponential, we typically only need to solve the root node of the MILP reformulation described below.
We summarize several favorable cases that frequently arise. First, if the LPEC contains only bound and complementarity constraints, it can be solved in linear time (Kirches2022, , Propositions 2.3 and 2.4). This algorithm is incorporated into our implementation. Second, at a feasible point $x^{k}$, if ${\mathcal{I}}_{00}(x^{k})=\emptyset$, then the LPEC (12) reduces to a standard LP. According to (10), the LPEC can then be solved by enumerating all branch LPs, which is not computationally expensive when $|{\mathcal{I}}_{00}(x^{k})|$ is small. Third, if $x^{k}$ is not B-stationary, it is unnecessary to solve all branch LPs; it suffices to find a feasible direction $d$ with $\nabla f(x^{k})^{\top}d<0$ (cf. Corollary 4.2). Moreover, in the LPECs arising in Phase I of our method, it is sufficient to simply find a feasible point (cf. Theorem 4.1). In a MILP formulation, the solver can therefore be terminated as soon as such a feasible point is obtained. Section 5 shows that this results in a significant reduction in the number of LPs that need to be solved in the MILP algorithm. Finally, if $x^{k}$ is S-stationary, B-stationarity can be verified by solving a single LP (cf. Proposition 3.1). In summary, except when verifying B-stationarity, it is rarely necessary to solve the LPEC to global optimality – and even then, the task often reduces to solving a single LP.
In this paper, we focus on arguably the easiest-to-implement approach for LPECs, and which does not suffer from the drawbacks of regularization methods. We consider the equivalent MILP formulation of LPEC (15Hu2012 :
@@
$0\leq x^{k}_{1,i}\!+\!d_{1,i}\!\leq\!{y_{i}}M,$ $\forall i\in\{1,\ldots,m\},$ (18c)
$0\leq x^{k}_{2,i}\!+\!d_{2,i}\!\leq\!{(1-y_{i})}M,$ $\forall i\in\{1,\ldots,m\},$ (18d)
$\|d\|_{\infty}\leq\rho.$ (18e)
@@
The constraints (18c) and (18d) correspond to the big-M reformulation of the complementarity constraints in (15c), where $M$ is a sufficiently large positive constant. The binary variables $y\in\{0,1\}^{m}$ effectively select one of the branches of the L-shaped set defined by the complementarity constraints.
To reduce the problem complexity for a given $\rho$, if $d_{i}\leq\rho<x_{1,i}$ or $d_{i}\leq\rho<x_{2,i}$, we can fix the corresponding binary variable $y_{i}$ for every $i$ in a preprocessing step, as the other branch becomes unreachable. In the special case of selecting $\rho$ as in (16), all binary variables $y_{i}$ for $i\notin{\mathcal{I}}_{00}(x^{k})$ can be determined in advance, and it is observed that this contributes to the computational efficiency of our approach.
The MILP (18) can be solved with state-of-the-art commercial and open-source solvers such as Gurobi Gurobi or HiGHS Huangfu2018 . We discuss some special properties of the MILP (18), which are favorable for MILP solvers Gurobi ; Huangfu2018 .
@@
We denote the optimal objective value of (18) by $V_{\mathrm{MILP}}(x^{k})$, and of its LP relaxation, that is, a problem where $y\in\{0,1\}^{m}$ is replaced by $y\in[0,1]^{m}$, by $V_{\mathrm{LP}}(x^{k})$. Thus, it holds that $V_{\mathrm{LP}}(x^{k})\leq V_{\mathrm{MILP}}(x^{k})$. The tighter the relaxation, i.e., the smaller $V_{\mathrm{MILP}}(x^{k})-V_{\mathrm{LP}}(x^{k})$ is, the less branching may be needed in MILP solvers Wolsey1998 . Big-M formulations can be inefficient because they can lead to rather loose relaxations of the MILP Hu2008 . However, for $\rho<\frac{M}{2}$, the trust-region constraint (18e) always dominates the bounds on $d$ from the big-M constraints (18c) and (18d). Thus, $M$ only needs to be large enough to ensure the feasibility of (18c)–(18d). A suitable choice is $M=\max(x_{1}^{k},x_{2}^{k})+\rho$. In addition, observe that the constraint (18e) makes the feasible set of the MILP compact. Moreover, for feasible points $x^{k}$, the trust-region radius $\rho$ can be arbitrarily small, leading to a smaller $V_{\mathrm{LP}}(x^{k})$ and thus tighter relaxations of the MILP.
Furthermore, given a feasible complementarity pair, e.g. from the previous BNLP, a feasible binary vector $y$ can be constructed, providing an upper bound for $V_{\mathrm{MILP}}(x^{k})$. The availability of an upper bound can reduce branching in branch-and-bound MILP algorithms Wolsey1998 , and our implementation includes the procedure for computing such a feasible binary vector.
Our numerical experiments indicate that these factors typically result in not large numbers of LP that need to be solved when solving the MILP (18). Interestingly, in our experiments the MILP formulations for LPECs are often substantially faster than solving the LPEC (15) with noncombinatorial regularization-based methods.
We discuss a special case that frequently arises in our numerical experiments when $x^{*}\in\Omega$ is an S-stationary point. In this case, the MILPs are particularly easy to solve since the relaxation is tight, i.e., $V_{\mathrm{MILP}}(x^{*})=V_{\mathrm{LP}}(x^{*})$. Thus, verifying B-stationarity reduces to solving a single LP. We prove this for the reduced LPEC (12). However, if $\rho<\bar{\rho}$ (with $\bar{\rho}$ from (16)) and the preprocessing step for fixing integers is applied, then the feasible sets of the full MILP (18) and reduced MILP (19) coincide.
Proposition 3.1
Let $x^{*}\in\Omega$ be an S-stationary point of the MPEC (1)
at which the MPEC-ACQ holds. Consider the MILP reformulation of the reduced LPEC (12):
$\underset{d\in{\mathbb{R}}^{n}\!,y\in\{\!0,1\!\}^{|}{\mathcal{I}}_{00}^{*}|}{\mathrm{min}}\;\quad$ $\nabla f(x^{*})^{\top}d$ (19a)
s.t. $c_{i}(x^{*})+\nabla c_{i}(x^{*})^{\top}d\geq 0,$ $\forall i\in{\mathcal{A}}(x^{*}),$ (19b)
$x_{1,i}^{*}+d_{1,i}=0,\ x_{2,i}^{*}+d_{2,i}\geq 0,$ $\forall i\in{\mathcal{I}}_{0+}(x^{*}),$ (19c)
$x_{1,i}^{*}+d_{1,i}\geq 0,\ x_{2,i}^{*}+d_{2,i}=0,$ $\forall i\in{\mathcal{I}}_{+0}(x^{*}),$ (19d)
$0\leq x^{*}_{1,i}\!+\!d_{1,i}\!\leq\!{y_{i}}M,$ $\forall i\in{\mathcal{I}}_{00}(x^{*}),$ (19e)
$0\leq x^{*}_{2,i}\!+\!d_{2,i}\!\leq\!{(1-y_{i})}M,$ $\forall i\in{\mathcal{I}}_{00}(x^{*}),$ (19f)
$\|d\|_{\infty}\leq\rho.$ (19g)
For $\rho<\frac{M}{2}$, $d=0$ is a global minimizer of both the MILP (19) and its relaxation with $y_{i}\in[0,1]$.
Proof.
First we show that $d=0$ is global minimizer of the MILP (19). It follows from Theorem 2.1 that if $x^{*}\in\Omega$ is S-stationary it is also B-stationary. Hence, $d=0$ is a global minimizer of the LPEC (6) and of ${\mathrm{LPEC}}(x^{*},\rho)$ in (12), because the additional trust-region constraint in the latter remains inactive. The MILP (19) and LPEC (12), are equivalent, hence $d=0$ is a global minimizer of the MILP.
[…]
s.t. $c_{i}(x^{*})+\nabla c_{i}(x^{*})^{\top}d\geq 0,\ \forall i\in{\mathcal{A}}(x^{*}),$ (20b)
$x_{1,i}^{*}+d_{1,i}=0,\ x_{2,i}^{*}+d_{2,i}\geq 0,\;$ $\forall i\in\mathcal{I}_{0+}(x^{*}),$ (20c)
$x_{1,i}^{*}+d_{1,i}\geq 0,\ x_{2,i}^{*}+d_{2,i}=0,\;$ $\forall i\in\mathcal{I}_{+0}(x^{*}),$ (20d)
$x_{1,i}^{*}+d_{1,i}\geq 0,\ x_{2,i}^{*}+d_{2,i}\geq 0,\;$ $\forall i\in\mathcal{I}_{00}(x^{*}),$ (20e)
$\|d\|_{\infty}\leq\rho$ (20f)
[…]
In the relaxed MILP, $y_{i}\in\{0,1\}$ in (19) is replaced by $y_{i}\in[0,1]$. Next, from the upper bounds in (19e) and (19f), we have that $\frac{d_{1,i}}{M}\leq y_{i}\leq 1-\frac{d_{2,i}}{M}$. Since the MILP is feasible, it follows that this interval is nonempty if $\frac{d_{1,i}}{M}\leq 1-\frac{d_{2,i}}{M})$ holds. This implicitly defines the constraint: $d_{1,i}+d_{2,i}\leq M$. If we choose $\rho<\tfrac{M}{2}$, then the upper bounds in (19e) and (19f) are never active. For such a choice, the feasible sets of the relaxed MILP (19) and the LP (20) coincide for all components $d\in{\mathbb{R}}^{n}$, for all $y\in[0,1]^{|{\mathcal{I}}_{00}|}$. Since the objective of (19) does not depend on $y$, it follows that $d=0$ is also a global minimizer of the relaxation of (19). In conclusion, $d=0$ is a global solution of the MILP (19) and LPEC (12) (verifies B-stationarity), relaxed LP (verifies S-stationarity) and relaxed MILP. ∎
Remark 3.1 (On Algorithm 1 and mixed-integer algorithms)
Any MPEC (1) can be reformulated into a mixed-integer NLP (MINLP) using the big-M reformulation:
$\underset{x\in{\mathbb{R}}^{n},y\in\{0,1\}^{m}}{\mathrm{min}}\;\quad$ $f(x)$ (21a)
s.t. $c(x)\geq 0,$ (21b)
$0\leq x_{1,i}\leq y_{i}M,$ $\forall i\in\{1,\ldots,m\},$ (21c)
$0\leq x_{2,i}\leq(1-y_{i})M,$ $\forall i\in\{1,\ldots,m\}.$ (21d)
However, the convergence properties our method and MINLP methods differ in several aspects. Most MINLP methods converge to the global optimum when the relaxed integer problem is convex Belotti2013 , but Algorithm 1 does not. This distinction is illustrated by the MPEC example in Example 3.2, whose equivalent MINLP (after relaxing $y\in\{0,1\}^{m}$ to $y\in[0,1]^{m}$) yields a convex quadratic program. There, Algorithm 1 converges only to a B-stationary point, not necessarily the global optimum.
While Algorithm 1 solves NLPs and MILPs similar to many MINLP methods Belotti2013 , there are significant differences in termination criteria and the particular subproblems that are solved. Nonlinear branch-and-bound algorithms terminate when the optimality gap falls below a threshold Belotti2013 , which does not directly relate to B-stationarity unless the global optimum is found. Outer approximation algorithms also solve sequences of NLPs and MILPs Belotti2013 , but their MILPs include linearized constraints from multiple different linearization points, making them different from the MILP (18). Consequently, it is not immediately clear when these methods terminate at B-stationary points. In summary, Algorithm 1 is not a MINLP algorithm; rather, it employs MILP solvers as a computational tool for solving LPECs. It aims to find a locally stationary point of an MPEC, which—unlike in MINLPs – is a well-defined notion.
3.3 Phase I: Finding a feasible branch NLP
In this section, we introduce an algorithm to find an initial feasible BNLP and its corresponding solution $x^{0}$, or return a stationary point of an auxiliary problem that minimizes a measure of constraint infeasibility.
Regularization-based methods may not always find B-stationary points, but they can be efficient in finding (almost) feasible points. A Phase I method is proposed that combines solving LPECs with a regularization-based method Hoheisel2013 ; Kim2020 .
In principle, one can use any such method. In our exposition, we focus on Scholtes’ global relaxation method, which performs well in numerical experiments Kim2020 ; Hoheisel2013 ; Nurkanovic2024b . It solves a sequence of NLPs of the form:
$\underset{x\in{\mathbb{R}}^{n}}{\mathrm{min}}$ $f(x)$ (Reg($\tau$))
s.t. $c(x)\geq 0,$
$x_{1},x_{2}\geq 0,$
$x_{1,i}x_{2,i}\leq\tau,$ $\forall i\in\{1,\ldots,m\}.$
We compute $x^{*}(\tau)$, a stationarity point of Reg($\tau$), which is not necessarily feasible for the MPEC (1). An LPEC is then constructed at $x^{*}(\tau)$ (for some $\tau>0$), and its solution determines an active set corresponding to a feasible BNLP.
It is well-known that certifying infeasibility of a general nonconvex optimization problem is difficult. In many NLP solver implementations, if the globalization strategy cannot make sufficient progress toward a feasible point, a feasibility restoration phase is triggered, cf. (Fletcher2002, , Sec. 3.3.), (Waechter2006, , Sec 3.3.), and (Byrd2006, , Sec. 6). Thereby an auxiliary NLP – which minimizes the constraint violation – is solved Byrd2006 ; Fletcher2002 ; Waechter2006 . If its stationary point is infeasible for the original problem, the solver terminates. Infeasibility of the original problem is sufficient but not necessary for this outcome. This outcome is often called local infeasibility Waechter2006 , which is not a global certificate of infeasibility.
@@ 3.3 Phase I: Finding a feasible branch NLP
Regard the MPEC (1) and the corresponding $\mathrm{Reg}(\tau)$ in  Reg($\tau$) for some $\tau>0$. If $\mathrm{Reg}(\tau)$ is infeasible then the same holds for the MPEC (1).
@@
On the other hand, if Reg($\tau$) is feasible for all $\tau$, then also the MPEC (1) is feasible as well. We expect that if $\tau^{k}$ is sufficiently small, then the feasible set of $\mathrm{LPEC}(x^{*}(\tau^{k}),\rho)$ is locally a good approximation of the original feasible set, and can help identifying a feasible BNLP. Algorithm 2 details such a procedure. Before going into details, we illustrate how an LPEC solution may fail or succeed in identifying a feasible BNLP for different values of $\tau$.
@@
[…]
with $a>1$. In the example, we set $a=1.1$. For larger values of $\tau$ both branches are feasible for the LPEC (15), but only the branch $x_{1}=0,x_{2}\geq 0$ is feasible for the MPEC. On the other hand, for smaller values of $\tau$, also in the LPEC, only the branch $d_{1}=0,\geq d_{2}\geq 0$ is feasible, thus the LPEC predicts only active set $({\mathcal{I}}_{1}(x^{*}(\tau)+d),{\mathcal{I}}_{2}(x^{*}(\tau)+d))$, which results in a feasible BNLP. Fig. 3 illustrates two sample solutions for $\mathrm{Reg}(\tau)$ and the corresponding $\mathrm{LPEC}(x^{*}(\tau),\rho)$. In this example, the smaller the parameter $a$ is the smaller $\tau$ must become.
@@
Figure 3: Illustrating of Example 3.3, where the LPEC (15) can correctly and wrongly identify a feasible BNLP depending on the linearization point $x^{*}(\tau)$. Top: the linearization point $x^{*}(\tau)$ is not close enough to the feasible set, so the LPEC admits an MPEC-infeasible branch. Bottom: $x^{*}(\tau)$ sufficiently close, so all LPEC branches are MPEC-feasible.